Steve recently inherited $70,000 that he feels he can invest and earn 4% compounded semiannually. If this is the only investment he can make toward his goal, how many years will it take before Steve has $100,000.

a. 18 years
b. 4.5 years
c. 6 years
d. 9 years
Answer #1

Ans d. 9 years

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV = PV (1 + r )n
100000= 70000*(1+4%/2)^n
(1+2%)^n = 100000/70000
(1+2%)^n = 1.4285714
n = 18
Ans 18 periods = 9 years